# Corner modes in Non-Hermitian long-range model

### v1

@@ IV Summary and Conclusion
To summarize, in this article, we consider a long-range hopping model in the presence of non-Hermiticity to investigate the emergence of second-order topological phases. By exploring the real part of the complex energy spectrum for the first and second NN NH models under OBC, we find that the former model only hosts four zero-energy corner modes while the latter model can host four as well as sixteen zero-energy corner modes as the hallmark of the NH SOTI phases. We compute the real space invariant namely, bi-orthogonalized QWN, by keeping in mind the non-Bloch form of the momentum to uniquely characterize the different topological phases. The phase boundaries, captured by the above invariant, can successfully mimic the emergence of NH SOTI phases out of the trivial phases, as demonstrated by the complex energy dispersion under OBC for both the first and second NN models. The topological phase boundary between two different topological phases, observed in the second NN model, can be anticipated from the complex energy spectrum under PBC for the above model. In the future, one can include disorder to study the exceptional topological Anderson insulators hosting corner modes where a generalized version of the presently adopted real space topological index will have to be examined.

### v2

@@ IV Summary and Conclusion
To summarize, in this article, we consider a second NN hopping model in the presence of non-Hermiticity to investigate the emergence of second-order topological phases. By exploring the real part of the complex energy spectrum for the first and second NN NH models under OBC, we find that the former model only hosts four zero-energy corner modes while the latter model can host four as well as sixteen zero-energy corner modes as the hallmark of the NH SOTI phases. We compute the real space invariant, namely, bi-orthogonalized QWN, by keeping in mind the non-Bloch form of the momentum to uniquely characterize the different topological phases. The phase boundaries captured by the above invariant can successfully mimic the emergence of NH SOTI phases out of the trivial phases, as demonstrated by the complex energy dispersion under OBC for both the first and second NN models. The topological phase boundary between two different topological phases, observed in the second NN model, can be anticipated from the complex energy spectrum under PBC for the above model. In the future, one can include disorder to study the exceptional topological Anderson insulators hosting corner modes where a generalized version of the presently adopted real space topological index will have to be examined.
@@ Acknowledgments
Appendix A Spatial symmetries and localization of corner states for asymmetric $\gamma$
Figure 6: In 2D domain, we illustrate the LDOS spectrum associated with the $E=0$ states choosing different values of $\gamma_{x}$ and $\gamma_{y}$: (a) $\gamma_{x}=\gamma_{y}=0.6$, (b) $\gamma_{x}=-\gamma_{y}=0.6$, (c) $\gamma_{x}=-\gamma_{y}=-0.6$, (d) $\gamma_{x}=\gamma_{y}=-0.6$, (e) $\gamma_{x}=0.6,\gamma_{y}=0.3$, and (e) $\gamma_{x}=0.3,\gamma_{y}=0.6$. We choose $m_{0}=1.0$, while the other model parameters remain the same as mentioned in Fig. 1.
In Table 1, we list all the symmetries that the model Hamiltonian $H_{\gamma}(\boldsymbol{\mathrm{k}})$ breaks or preserves. As discussed in the main text, the mirror rotation symmetry $M_{xy}$ plays a crucial role in the localization of the corner states. In Fig. 6, we demonstrate the LDOS associated with the $E=0$ states choosing different values of $\gamma_{x}$ and $\gamma_{y}$. Note that, in the main text, we always consider $\gamma_{x}=\gamma_{y}=\gamma$. In Figs. 6(a-d), we illustrate the case when $\gamma_{x}$ and $\gamma_{y}$ carry the same amplitude but can have different signs for a NH first NN Hamiltonian model. We observe that depending upon the signs of $\gamma_{x}$ and $\gamma_{y}$, the corner modes occupy different corners of the system. In contrast, when $\gamma_{x}\neq\gamma_{y}$, i.e., the mirror rotation symmetries are broken, the corner states can occupy more than one corner [see Figs. 6(e,f)]. However, it is to be noted that the localization at different corners carry different weights.
Figure 7: In panels (a) and (b), we demonstrate ${\rm Re}[E]$ and ${\rm Abs}[E]$ close to the rightmost phase transition line in Figs. 3(d,f), choosing different system sizes respectively.
Appendix B Finite size scaling with second-NN NH hopping model
As discussed in the main text, the rightmost part of the phase boundary (in Figs. 3(d,f)) corresponding to the second NN NH Hamiltonian encounters finite-size scaling. Here, we explicitly exhibit variation of the eigenvalue spectra considering different system sizes. In particular, we demonstrate the real (${\rm Re}[E]$) and absolute part (${\rm Abs}[E]$) of the eigenvalue spectra close to the rightmost phase transition line of Figs. 3(d,f) choosing different system sizes ($N=12,16,20,24$) in Fig. 7. One can evidently observe from Figs. 7(a,b) that as we increase the system size, the zero-energy states in the eigenvalue spectra move towards the analytically obtained phase transition line (red). However, the translation towards the phase transition line as a function of the system size appears to be slower in nature.